OpenStudy (anonymous):

A 100-litre tank filled with water is leaking water at the rate of 2.5 litres per hour. a) How long will it take for the tank to be half full? Once it is half full, water is added to the tank at the rate of 3 litres per hour. b) How much longer will it take until the tank is full again?

OpenStudy (anonymous):

I've done and got part a correct correct already. I need help with part b.

OpenStudy (lgbasallote):

for it to be full it needs 50 litres...since it was half right

OpenStudy (lgbasallote):

so just do 50/3

OpenStudy (anonymous):

no but it is still leaking 2.5L per hour

OpenStudy (lgbasallote):

ohh well that's different

OpenStudy (anonymous):

so yes we need 50L more to be filled back to 100L(the starting point)

OpenStudy (anonymous):

and although we are adding 3L per hour 2.5L is still leaking per hour so only 0.5L is actually added.

OpenStudy (lgbasallote):

so it's like 0.5 litres is just added per hour

OpenStudy (lgbasallote):

however if you do 50/0.5 you wont get the right answer

OpenStudy (anonymous):

yep, that's why I need help.

OpenStudy (lgbasallote):

im thinking you neeed to do 50/0.5 THEN divide by 2

OpenStudy (lgbasallote):

because by dividing by 2 you're going to see when it first got full....my hypothesis only though

OpenStudy (ganpat):

leakage at a rate of 2.5 l/ hr.. so for 50 liters = 50/2.5 = 20 hrs.. Now to full the tank again to 100 liters, 50/2.5 + 50/ 3 = 36.67 hrs ... what say guys ??

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

The answer is actually 100hours for part b but I do not know how it is achieved.

OpenStudy (anonymous):

actually I got it.

OpenStudy (lgbasallote):

100? 50/0.5???

OpenStudy (anonymous):

yeh

OpenStudy (ganpat):

sorry i missed a step... leakage at a rate of 2.5 l/ hr.. so for 50 liters = 50/2.5 = 20 hrs.. so 1/2 tank is emptied in = 1/20 hrs.. Now, time taken to fill 50liters at a rate of 3 l/hr.. = 50/3 = 16.67 hrs.. So, 1/2 tank is filled in = 1/16.67 hrs.. Now when both happens together, = (1/16.67) - (1/20) = 3.33/ (20 * 16.67) = 1/100.12.. so tank will be filled in 100.12 hrs...

OpenStudy (lgbasallote):

that's wrong o.O well it's partially right...but that's not when it will reach full tank

OpenStudy (anonymous):

50/0.5 = 100 I believed I did 50 x 0.5 = 25 so I must have got it wrong by accidentally pressing the x and not the divide sign.

OpenStudy (lgbasallote):

i agree that in 100 hours the tank will reach fulll...but the tank will reach full before 100 hours too

OpenStudy (anonymous):

how?

OpenStudy (ganpat):

@lgbasallote : actually little more than 100 hrs... i have calculated that..

OpenStudy (lgbasallote):

hmmm

OpenStudy (lgbasallote):

nevermind...it's happening simultaneously

OpenStudy (anonymous):

@Ganpat, ur method is too complicated for my likes sorry.

OpenStudy (anonymous):

and although ur method seems to work the answer I've got of exactly 100 seems to be more correct.

ganeshie8 (ganeshie8):

you can use distance = time * (effective speed)

OpenStudy (ganpat):

@JayDS : 50/0.5 , might be the right answer ion this case.. but it may not be true for different problems...

ganeshie8 (ganeshie8):

here, distance is related to - water accumalted speed is related to - flow rate

OpenStudy (anonymous):

lol but there's no distance here....

OpenStudy (ganpat):

@JayDS : Its not actually.. its a trick to solve this leakage problems.. always convert them into per unit... calculations will be easy... and it always works..

OpenStudy (anonymous):

and we do not need to find distance.

OpenStudy (anonymous):

ok, I'm not sure if it works using @ganeshie8's way but thanks.

ganeshie8 (ganeshie8):

it works. its a generic relation. it works for any speed/distance/rate situations

OpenStudy (anonymous):

okok, it seems to relate to Physics which I do not learn lol, so I rather keep it simple as possible.

ganeshie8 (ganeshie8):

its not physics. its just a relation. forget about the terminology i used if it confuses

OpenStudy (anonymous):

and @ganpat I'll give u the medal since u tried so hard and to help me, u troll. lol

OpenStudy (ganpat):

ha ha, .. @JayDS: but its a really good, simple and prominent method.. SO try to grasp it.. i can help u understand...

ganeshie8 (ganeshie8):

yes.. and, if the relation quoted above seems new, here is another way : 1) amount of water coming into tank = 3*t 2) amount of water leaving out of tank = 2.5*t 3) amount of water in tank = 3t-2.5t 50 = 3t-2.5t 50 = .5t t = 100